two shuffleboard of equal mass,one orange and the other yellow,are involved in a elastic,glancing collision.the yellow disk is initially at rest and is struck by the orange disk moving with a speed of 5.00 m/s.after collision,the orange disk moves along a direction that makes an angle of 37.0 with its initial direction of motion.the velocities of the two disks are perpendicular after the collusion.determine the final speed of each disk


1
Expert's answer
2021-10-08T09:30:16-0400

Solution.

Before collision:

"mv1 + mv2 = const"

v2 = 0, so

"mv1 = const"

After collision:

Horizontal direction

"mv1 = mv^{'}1cos(37^o)+mv^{'}2sin(90^o-53^o)"

"v1 = v^{'}1cos(37^o)+v^{'}2sin(90^o-53^o)"

Vertical direction

"v^{'}1sin(37^o)-v^{'}2sin(53^o) = 0"

"v^{'}1 = v^{'}2 \frac{sin(53^o)}{sin(37^o)}"

"v1 = v^{'}2 \frac{sin(53^o)}{sin(37^o)}cos(37^o)+v^{'}2cos(53^o)"

"v1sin(37^o) = v^{'}2"

"v2' = 5 \times sin(37^o) = 3.01 m/s"

"v^{'}1 = 3.01 \frac{sin(53^o)}{sin(37^o)} = 3.99 m/s"

Answer:

v(orange) = 3.99 m/s

v(yellow) = 3.01 m/s